# Heavy-element damage seeding in proteins under XFEL illumination

### v4

@@ VII Conclusions
For the most part, this work has restricted analysis to targets where heavy elements make up 1% of the atomic population, but this is far from sufficient to generalize the influence of heavy atoms across the varied ratios seen in real targets, including targets containing multiple heavy elements. However, it is evident that experimental differences generally considered marginal in traditional crystallography can substantially affect the amount of radiation damage suffered by targets in the ultrafast regime. This complexity emphasizes the value of using theoretical modeling to inform SFX experimental design – a role it is already fulfilling [36] – particularly as a tool for gauging the viability of successful refinement in the high-intensity regime. The zero-dimensional framework employed in AC4DC can capably examine the complete electronic damage dynamics across a large number of candidate pulse parameterizations without significant investment of computational resources. For studies concerned with the ions’ motions, the simulation may be integrated within a hybrid plasma-MD framework [33, 56], where delegation of the ultrafast electron dynamics to a zero-dimensional model makes simulating the molecular dynamics of 10–100 nm scale structures feasible.
@@ Acknowledgements
We thank A/Prof. Nadia Zatsepin, Swinburne University of Technology, Hawthorn, Australia, for feedback on the paper’s discussions of SFX experimental methods and Thippayawis Cheunchitra, University of Melbourne, Parkville, Australia, for inspiring examination of the influence of absorption edges on the damage landscape.
@@ Appendix A Numerical Method
Simulating scattering off a protein from an XFEL pulse consists of two stages -
- 1.
2.
Perform a spatially-resolved Monte Carlo simulation of the elastic scattering pattern, with the ionic states of the atoms in the structure sampled from at each discrete timestep.
A.1 Scattering pattern
The scattering pattern is calculated as [2, 30]
Here, is the momentum transfer corresponding to pixel , is the fluence of the (unpolarized) incident beam, and is the classical electron radius. is defined by approximating each atom as remaining fixed in its initial position , with each form factor (the Fourier transform of the atom’s electron density) corresponding to a state with probability , giving
A.2 Cubic spline representation of the free electrons
The collision kernel on the right hand side of the Boltzmann equation, Equation (1), takes the form of a sum over distinct processes,
The solution of these coupled rate equations presents serious numerical challenges:
@@ A.2 Cubic spline representation of the free electrons
The three-body recombination term is quadratic in the free electron distribution , leading to worst-case complexity if the representation of is dimensional.
@@
The electron-electron recombination term depends on derivatives of the free electron distribution.
@@
The standard approach [84] to non-LTE plasma simulation solves the Boltzmann equation using finite differencing of . We take a more general approach, expanding with respect to a time-invariant basis , contracted with time-varying expansion coefficients and multiplied by an explicit weight function ,
@@
The first-order electron density function has two components – , the continuum energy distribution of the free continuum, and , the time-dependent density of ions in atomic configuration .
We separated the time-dependent energy distribution of free electrons from the discrete probability distribution capturing the classical populations of each atomic state . The Boltzmann and master equations (1) and (2) then generate a deterministic time evolution of the electrons’ classical energy-state distribution.
Further numerical methods employed included: (I) An asynchronous implicit-explicit (IMEX) method, stepping the stiff but computationally cheap free electron interactions with much shorter steps than the bound-bound and bound-free contributions. (II) Adaptive time steps to avoid divergence in .
A.3 Explicit form of numerical couplings
@@
A.3.1 Electron-Electron interactions
@@
A.3.2 Photoionization and Auger decay
@@
A.3.3 bound-free interactions
@@
A.4 Dynamic grid implementation
The plasma code’s cubic spline treatment has previously been shown to well-approximate the evolution of relatively static Boltzmann-governed systems with as few as 10 energy grid points/knots [86]. However, this requirement grew by an order of magnitude when the initial distribution of particles was modestly (relative to the energy scale of the present work) displaced from equilibrium with a thermal bath. The unbound electron continuum of a biological target under XFEL illumination sees an acute difference between its state at early times and its partially equilibrated state near the end of the pulse. Moreover, these systems see two complications not present in the systems considered by Khurana et al.: (I) The narrow, low-energy peak of the Maxwellian at early times is difficult to fit due to the rigid energy conservation condition of . (II) The thermal bath is substituted for sharp, high-energy emission profiles, which, like the Maxwellian, shift significantly over the course of the pulse as they relax through electron interactions [34].
An adaptive grid was implemented to address these issues. A set of static low-density regions spans the full energy range and a set of dynamic high-density regions spans (I) the Maxwellian and (II) up to 4 of the most dominant high-energy peaks. We define a set of energy ranges (regions), each with an associated knot density function that is only non-zero within the region. The local knot density of the full grid is then defined as
The high-density regions of the electron continuum are dynamically updated throughout the simulation to follow their respective features. A high density of knots supports the sharp features in the continuum at early times, then redistribute to continue to support the growing and shifting peaks as the electron population equilibrates. Accurate fitting of the continuum at early times appears to be particularly crucial, so a partial run of the simulation with a ‘guess grid’ is used to obtain the initial grid for the actual simulation. This flexible approach drastically reduced the number of knots necessary to achieve convergence in the evolution of the ionic states.
@@ A.4 Dynamic grid implementation
Appendix B
The standard theoretical measure of electronic damage [2] is
where is the momentum transfer corresponding to a pixel (in some other studies, the Bragg peak), and and are the normalized irradiances scattered by the system in the cases where electronic damage is, respectively, ignored or included.
We note that the ‘rule of thumb’ that 0.15–0.20 implies a recoverable structure [2, 29, 20, 32] is difficult to reconcile with certain experimental concerns. The essence of this matter may be distilled into two points. (I) The original introduction of a notional cutoff was predicated on being an achievable limit for the experimental measure of deviation within the merged dataset () [2]. However, data quality factors are not predictive of refinement quality [87, 88]. (II) measures only damage-induced noise, while experimental factors account for all sources of noise. Both and the standard refinement measure are mostly confined to a value range of across macromolecules deposited in the Protein Data Bank [89, 2], so it is ill justified to discount a single source of noise which in isolation produces an factor on the same order–at least not without experimental support.
Even in the regime where radiation damage is minimal and would be near 0, an below 0.15 is rarely seen in refined models of macromolecules [89]. The confounding factors responsible for this difficulty may be magnified by radiation damage, making a cutoff based solely on trends in factors of deposited structures ill-defined. This suggests that a cutoff for that reliably gauges the limit of tolerable damage in SFX would need to be informed by (presently lacking) empirical data on how damage impacts .
@@
Appendix C Influence of cascade energy with controlled seeding rate
@@
Increasing the X-ray energy above an absorption edge decreases the photoabsorption coefficient, but also raises the energy of its seeded cascades. To isolate the relationship between a cascade’s energy and the damage it causes, AC4DC was modified to model the ionization of the light-atom control with additional electrons injected at a rate independent of their energy, under a Gaussian 10 keV ph-2, 15.0 fs FWHM pulse. The injection rate was fixed at three times the photoionization rate of the undamaged target for a given incident intensity, coarsely corresponding to the ionization rate of the Zn-doped lysozyme by its Zn ions under the same pulse conditions. (Zn was chosen due to showing a particularly distinct local maximum above its K-edge.) The resulting relationship between the electron source energy and the damage suffered by the target (Fig. 11) shows strong similarities to the relationship between LPE and damage in the Zn-doped target shown in Fig. 7(b). Notably, showing a local maximum at a similar position and magnitude.

### v5

@@ VI Conclusions
This work has restricted its broader analysis of the electronic damage landscape to targets where heavy elements make up 1% of the atomic population, but this is far from sufficient to generalize the influence of heavy atoms across the varied ratios seen in real targets, including targets containing multiple heavy elements. However, it is evident that experimental differences generally considered marginal in traditional crystallography can substantially affect the amount of radiation damage suffered by targets in the ultrafast regime. This complexity emphasizes the value of using theoretical modeling to inform SFX experimental design – a role it is already fulfilling [37] – particularly as a tool for gauging the viability of successful refinement in the high-intensity regime. The zero-dimensional framework employed in AC4DC can capably examine the complete electronic damage dynamics across a large number of candidate pulse parameterizations without significant investment of computational resources. For studies concerned with the ions’ motions, the simulation may be integrated within a hybrid plasma-MD framework [34, 58], where delegation of the ultrafast electron dynamics to a zero-dimensional model makes simulating the molecular dynamics of 10–100 nm scale structures feasible.
@@ Acknowledgements
We thank A/Prof. Nadia Zatsepin, Swinburne University of Technology, Hawthorn, Australia, for feedback on the discussions of SFX experimental methods and Thippayawis Cheunchitra, University of Melbourne, Parkville, Australia, for inspiring examination of the damage landscape above deep absorption edges.
Data availability
Simulation input files for Sec. III may be found at https://github.com/Phoelionix/AC4DC/tree/master/input/lysozyme. Input files for Sec. IV may be found at https://github.com/Phoelionix/AC4DC/tree/master/input/templates/doped_targets. Additional data used in this work is available upon reasonable request to the corresponding author.
@@ Appendix A Numerical method
A.1 Cubic spline representation of the free-electron energy distribution
The collision kernel on the right-hand side of the Boltzmann equation, Equation (1), takes the form of a sum over distinct processes,
The solution of these coupled rate equations presents several numerical challenges:
@@ A.1 Cubic spline representation of the free-electron energy distribution
The three-body recombination term is quadratic in the free-electron energy distribution , leading to worst-case complexity if the representation of is dimensional.
@@
The electron-electron recombination term depends on derivatives of .
@@
The standard approach [87] to non-Maxwellian plasma simulation solves the Boltzmann equation using finite differencing of . We take a more general approach, expanding with respect to a time-invariant basis , contracted with time-varying expansion coefficients and multiplied by an explicit weight function ,
@@
We separated the time-dependent energy distribution of free electrons from the discrete probability distribution capturing the classical populations of each atomic state . The Boltzmann and master equations (1) and (2) then generate a deterministic time evolution for the classical energy-state distribution of the electrons.
Further numerical methods employed included: (I) An asynchronous implicit-explicit (IMEX) method, stepping the stiff but computationally cheap free-electron interactions with much shorter steps than the bound-bound and bound-free contributions. (II) Adaptive time steps to avoid divergence in .
A.2 Explicit form of numerical couplings
@@
A.2.1 Electron-Electron interactions
@@
A.2.2 Photoionization and Auger decay
@@
A.2.3 bound-free interactions
@@
A.3 Dynamic grid implementation
The cubic spline treatment of the plasma code has previously been shown to well-approximate the evolution of relatively static Boltzmann-governed systems with as few as 10 energy grid points/knots [89]. However, this requirement grew by an order of magnitude when the initial distribution of particles was modestly (relative to the energy scale of the present work) displaced from equilibrium with a thermal bath. The unbound electron continuum of a biological target under XFEL illumination sees an acute difference between its state at early times and its partially equilibrated state near the end of the pulse. Moreover, these systems see two complications not present in the systems considered by Khurana et al.: (I) The narrow, low-energy peak of the Maxwellian at early times is difficult to fit due to the rigid energy conservation condition of . (II) The thermal bath is substituted for sharp, high-energy emission profiles, which, like the Maxwellian, shift significantly over the course of the pulse as they relax through electron interactions [35].
An adaptive grid was implemented to address these issues. A set of static low-density regions spans the full energy range and a set of dynamic high-density regions spans i) the Maxwellian and ii) up to four of the most dominant high-energy peaks. We define a set of energy ranges (regions), each with an associated knot density function that is only non-zero within the region. The local knot density of the full grid is then defined as
The high-density regions of the electron continuum are dynamically updated throughout the simulation to follow their respective features. A high density of knots supports the sharp features in the continuum at early times, then redistribute to continue to support the growing and shifting peaks as the electron population equilibrates. Testing showed accurate fitting of the continuum at early times to be particularly crucial, so a partial run of the simulation with a ‘guess grid’ is used to obtain the initial grid for the actual simulation. This flexible approach drastically reduced the number of knots necessary to achieve convergence in the evolution of the ionic states.
@@ A.3 Dynamic grid implementation
Appendix B Influence of cascade energy with controlled seeding rate
@@
Increasing the X-ray energy further above an absorption edge raises the energy of the seeded EII cascades. At the same time, the number of these cascades decreases due to the reduced absorption cross-section at higher energies. To isolate the relationship between the energy of a primary electron and the damage caused by the resulting cascade, AC4DC was modified to model the ionization of the light-atom control with additional electrons injected at a rate independent of their energy. Each system was simulated under a 15.0 fs FWHM Gaussian pulse with a fluence of 10 keV ph-2. The injection rate was fixed at three times the photoionization rate of the undamaged target for a given incident intensity, coarsely corresponding to the ionization rate of the Zn-doped target by its Zn ions under the same pulse conditions. (Zn was chosen due to the distinct local maximum it produces, as shown in Fig. 6.) The resulting relationship between the electron source energy and carbon ionization (Fig. 10) shows strong similarities to the relationship between LPE and carbon ionization in the Zn-doped target shown in Fig. 6(b), including a similar position and magnitude for the local maximum.
Appendix C Model comparisons
C.1 Amorphous carbon – ddcMD
Ref. [35] used the MD code ddcMD to simulate the ionization of amorphous carbon of density 1 g cm-3 under a 2 mJ, keV, fs square pulse focused to a 100 nm FWHM focal size.Modeling with AC4DC was performedf taking a homogeneous intensity equal to the average in the 100 nm focus. This corresponds to a photon fluence of ph -2, an order of magnitude above the maximum fluence considered elsewhere in this work. AC4DC saw excellent agreement with ddcMD in both the evolution of the carbon ionic states and the free-electron energy distribution. Fig. 11 shows the Maxwellian fits to the thermal electrons at several snapshots in time predicted by ddcmD, overlaid on the corresponding normalized electron energy density distributions predicted by AC4DC.
C.2 Aluminum sheet – PICLS
A system corresponding to an infinite plane thick aluminum foil sheet was simulated by Ref. [59] using the particle-in-cell code PICLS under varying conditions. The code tracks individual electrons in space, and determines atomic transitions and impact ionization using Monte Carlo modeling. TBR is incorporated through enforcing an equation of state, which, as highlighted by Ref. [59], is less accurate than modeling TBR cross-sections. Unlike AC4DC, PICLS models ionization potential depression (IPD). IPD is a stronger effect in denser matter, and so would be expected to play a greater role in solid Al than in biological matter. As the EII rate is increased by IPD, it was expected that AC4DC would underestimate the rate of secondary ionization.
We consider two scenarios from this study corresponding to quite different pulse parameters. First, a pulse of 1.70 keV photons with a 80 fs FWHM Gaussian profile and peak intensity of W cm-2. Second, a pulse of 10.0 keV photons with a 20 fs FWHM Gaussian profile and peak intensity of W cm-2. Note that for the first pulse, the photoelectrons are ejected with a very low energy ( eV), so the cascades seeded by the 1.4 keV Auger electrons dominate the secondary ionization dynamics.
For the 1.7 keV pulse, the predictions for the evolution of the ion fractions by AC4DC and PICLS, shown in Fig. 12(a), see relatively good agreement, similar to that observed in the considered study with another collisional-radiative code, SCFLY. In comparison to PICLS, the ionization rate predicted by AC4DC is overall slightly lower. Inspection of the electron density distributions shown in Fig. 12(b) indicates that the electrons thermalize more slowly in the AC4DC simulation. This is consistent with a higher rate of collisional-ionization in the PICLS model. The different level of thermalization between the models limits the physical meaning of a comparison between temperatures. Similar can be said of the electron density distributions for the 10 keV pulse.
For the 10 keV pulse, the models again make similar predictions for the evolution of the free-electron energy distribution, as can be observed in Fig. 13. There is a surprisingly good match to the temperature of the thermal electrons at fs. However, this is again a limited comparison as the low-energy electrons have not completely relaxed to a Maxwellian in the AC4DC simulation, while they are nearly completely thermalized in the PICLS model. The low-energy electrons remained unrelaxed at fs in the AC4DC simulation (with a best-fit Maxwellian of 213.5 eV).
C.3 Glycine crystal – XMDYN
Ref. [65] used the Monte Carlo MD code XMDYN to model radiation damage in a glycine (H5C2NO2) crystal. The study modeled pulses of 10 keV photons with 10 fs FWHM Gaussian temporal profiles, at four different peak intensities, between – W cm-2. The simulation volume contained 1050 atoms, under periodic boundary conditions. The results of these simulations are shown in Fig. 14. The evolution of charges by element in equivalent AC4DC simulations, also shown in Fig. 14, was in good agreement with this modeling for the two higher fluence cases. In the two lower fluence cases, XMDYN predicts a higher level of ionization.
[H]
C.4 Silicon crystal – non-Maxwellian plasma code
We modeled the conditions a silicon crystal was subjected to in simulations performed by Ref. [55]. The considered study used an atomic code that was shown to provide good agreement with the HF method used by AC4DC. The physical model for the plasma dynamics is very similar to AC4DC, though the numerical approaches differ significantly. AC4DC was used to model the 4 keV and 8 keV pulses considered in this study, corresponding to energies near and from the 1.8 keV Si K-edge. AC4DC predicted a much higher rate of ionization and differing electron density continuums. Results for the 8 keV pulse are shown in Fig. 15.
C.5 Lysozyme.Gd – RADDOSE-XFEL
RADDOSE-XFEL [37] is a module of RADDOSE-3D [90] designed to measure the dose absorbed in XFEL experiments by tracking collisional ionization cascades. Unlike AC4DC, the code does not model high-density, strong-ionization effects such as TBR and the depletion of atomic orbitals. Since primary ionization drives the damage-seeding effect described in this study, the code was expected to reflect an outsized impact by heavy elements on the absorbed dose.
Table 1 summarizes the damage predictions of RADDOSE-XFEL and RADDOSE-3D applied to targets equivalent to the lysozyme.Gd target and its toy variants described in Sec. III. Values are scaled based off a sample of individually simulated photons, as described in Ref. [37]. The doses predicted by RADDOSE-3D are consistently higher than predicted by RADDOSE-XFEL because RADDOSE-3D assumes the energy of photons are instantaneously absorbed after photoionization [37]. The influence of heavy atoms is underestimated by this approximation; adding sulfur (light control to lysyozyme, water) and salt (lysozyme.Gd, 0.1 M Gd to lysozyme.Gd, 0.1 M Gd 10% NaCl) magnifies the dose by a much larger factor when the more accurate RADDOSE-XFEL model is used. Notably, accounting for sulfur increases the absorbed dose and average number of non-H ionizations by 49% and 45%, respectively. This effect is stronger than predicted by the AC4DC simulations, likely because RADDOSE-XFEL does not model the ionic states and so does not capture the suppression of EII that occurs as the target becomes increasingly ionized.
RADDOSE-XFEL predicts that accounting for Gd (lysozyme, water to lysozyme.Gd, 0.1 M Gd) does not increase ionization, but increases the dose by an amount comparable to that predicted by RADDOSE-3D. This result should be taken as unphysical. RADDOSE-XFEL models instantaneous energy deposition similar to RADDOSE-3D for many ionization processes of Gd. Indeed, Fig. 4 indicates the contribution of cascades seeded by Gd to global ionization should be similar to those seeded by Cl in the 10% NaCl solvent.
| Protein | Solvent (35.1% v/v) |
| Light-atom control, | water | 1931.4 | 325.7 | 1.278 | ||||||
| Lysozyme, | water | 2338.8 | 483.6 | 1.856 | ||||||
| Lysozyme.Gd, | 0.1 M Gd | 2923.2 | 1088.6* | 1.859* | ||||||
| Lysozyme.Gd, | 0.1 M Gd 10% NaCl | 3476.6 | 1304.8* | 2.779* |
RADDOSE-XFEL assumes instantaneous dose absorption for some Gd ionization processes.